The sum of 2 consecutive odd integers is 92. find the numbers.
Answer by bucky(2189)

Let x be the first odd integer. The next odd integer is two away from the first odd integer
because an even number appears between them.
.
Since x is the first odd integer, then next odd integer above it is x + 2. The problem
then says that if you add these two integers the result is 92.
.
In equation form this is:
.
x + (x+2) = 92
.
adding the x terms changes the left side to:
.
2x + 2 = 92
.
Eliminate the 2 on the left side by subtracting 2 from the left side. But subtracting
2 from the left side means you also have to subtract 2 from the right side. The result
of these subtractions is:
.
2x = 90
.
Solve for x by dividing both sides by 2 to get:
.
x = 90/2 = 45
.
Since 45 is the first odd number, the next odd number is 45 + 2 or 47. The answers
are 45 and 47.
.
Check by adding these two odd integers and you will see that they total 92, just as
specified by the problem.
.
Hope this helps you to understand how to identify odd integers as being 2 units above the
the preceding odd integer.